The heights of ten children selected at random from a given locality had a mean 63.2cm and variance 6.25cm. Test at 5% level of significance, the hypothesis that the children of the given locality are on average less than 65 cm. given for 9 degrees of freedom P(t>1.83)=0.05


1
Expert's answer
2021-07-16T01:33:21-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\geq65"

"H_1:\mu<65"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.05," "df=n-1=10-1=9" degrees of freedom, and the critical value for a left-tailed test is "t_c=-1.833113."

The rejection region for this left-tailed test is "R=\{t:t<-1.833113\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{63.2-65}{6.25/\sqrt{10}}=-0.910736"

Since it is observed that "t=-0.910736>-1.833113=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value for left-tailed, "\alpha=0.05, df=9," "t=-0.940736" is "p=0.193088," and since "p=0.193088>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is less than "65," at the "\alpha=0.05" significance level.